Students in a chemistry class convince their teacher to use the following "group grading" scenario. Students will all take the exam on their own; however, the grade they receive will be the mean of their test score with 4 other randomly selected classmates. Assume that the test scores for this particular exam are normally distributed with a mean of 74 and a standard deviation of 12 points. What is the probability that the mean test score for your group falls above an 80?
Answer #1

Solution :

Given ,

mean = = 74

standard deviation = = 12

P(x > 80) = 1 - P(x<80 )

= 1 - P[(x -) / < (80 - 74) /12 ]

= 1 - P(z < 0.5)

Using z table

= 1 - 0.6915

= 0.3085